Finco is a new firm that just paid an annual dividend of $1 a share. The firm plans to increase its dividend by 20% per year for the next four years and then decrease the growth rate to 5% annually. If the required rate of return is 10%, what is one share of this stock worth today?

$35.77

$34.77

$32.07

$48.59

$43.60
Answer #1

of value during abnormal Calculation stage pv@ 1000 - Dividends 1X120°. = 102 1-2x 120%. = 1-44 1-44x 120%. = 1.728 1.728 x 1Explaination. The value of stock taday is equal to present value of al future Cashflows expected from the stock discounted at